# Training-time Neuron Alignment through Permutation Subspace for Improving Linear Mode Connectivity and Model Fusion

### v1

@@ 3.2.1 Method Formulation
where $\odot$ denotes the element-wise Hadamard product, $\eta$ refers to the learning rate, and $\mathbf{g}_{i}$ is its gradients of the optimizer, such as SGD or Adam. After training for $E$ epochs, we validate the LMC with respect to the loss or accuracy barriers in Definitions 2.1 and 2.2. The method is notated as TNA-PFN.
@@
Discussions on differences between TNA-PFN and previous gradient-mask-based methods. Applying gradient masks is discovered in previous gradient compression literature of distribution optimization, but our method is different from the previous works in the aspects as follows. Motivation difference: Gradient compression is proposed for communication efficiency of distributed optimization while we study the training-time neuron alignment problem in LMC. Implementation difference: Gradient compression uses different random top-k gradient masks at each worker and changes the mask per communication iteration (Lin et al., 2018; Vogels et al., 2019); whereas, TNA-PFN uses the same random gradient masks at each model, fixes the mask, and independently trains the models without any communications; and FedPFN/FedPNU (presented in subsection 4.2) uses the same masks at each client’s local training and changes the mask per global communication round. Effect difference: Since the masks of workers are different and changing, previous gradient compression methods cannot learn in a consistent subspace of parameters, while we learn in a subspace by the same gradient mask so that some neuron weights are not updated.
3.2.2 Theoretical Analysis
We make a theoretical analysis about how TNA-PFN can improve LMC, shown in Theorem 3.2 (proof is in Appendix B). The main idea is to treat the linear interpolated landscape of the barrier as a function of parameter $\alpha$, and the connectivity can be depicted by the first and second derivatives of the function.
Theorem 3.2.
[…]
@@ 3.2.2 Theoretical Analysis
Table 1: The performances of post-matching methods after TNA-PFN. Interpolated Accuracy (Interp. Acc.) means the accuracy of the linearly interpolated model, i.e., $\mathcal{A}(0.5\mathbf{w}_{1}+0.5\mathbf{w}_{2})$. “Iter.” refers to the number of iterations in the post-matching methods, reflecting the computation costs.
@@
Theorem 3.2 shows how TNA-PFN improves LMC by masking some weights from updating (setting $\rho_{\bm{v}}$ and $\rho_{\bm{U}}$ larger than zeros), where the networks learn in a unified permutation subspace.
3.2.3 Empirical Results on Linear Mode Connecitivity
@@
Different model depths, widths, and architectures. In Figure 3, we conduct experiments on MLP with different hidden layers and widths. For MNIST (LeCun & Cortes, 2010), we find shallower and wider networks will not cause barriers, which is consistent with the previous observations (Entezari et al., 2022). For CIFAR-10 (Krizhevsky et al., 2009), the barriers always exist under various depths and widths. Our proposed TNA-PFN can obviously reduce the accuracy barriers from 0.3-0.4 to 0.1, and we also visualize the loss landscapes, which illustrate the barrier reductions. To see more results and illustrations, please refer to subsection C.1.
We study the LMC of simple CNN and ResNets and present the results in Figure 4. ResNets (He et al., 2016) have higher barriers than simple CNN, and the barriers are exacerbated when the networks are deeper or wider. It is suggested that TNA-PFN can lower the barriers under different architectures.
Generally, we observe that TNA-PFN has more dominant advantages when the models are wider, and the observations are: (1) the second figure in Figure 3: for CIFAR-10 when width increases, the barriers of vanilla training go up while the barriers of TNA-PFN go down; (2) Figure 4: the barrier reduction of TNA-PFN is more obvious for WRN56 compared with ResNet56. This observation is consistent with previous literature about permutation symmetry and linear mode connectivity (Entezari et al., 2022).
Table 2: Linear mode connectivity of non-random initializations. Dataset: CIFAR-10. Semi-trained CNN: The initialization is first trained on a disjoint subset of CIFAR-10 for 0.5 epoch. Pretrained ResNet18: Pretrained on ImageNet.
Table 3: Linear mode connectivity of multi-model fusion. The number of models is 5.
@@
The role of post-hoc neuron alignment methods after training-time alignment. We consider simulated annealing (SA) (Entezari et al., 2022) and weight matching (WM) (Ainsworth et al., 2022) after TNA-PFN in Table 1. SA requires large computations, and we notice the improvements are also marginal. Under limited computation budgets (10 or 100 iterations), we find that TNA-PFN can reach a higher result than vanilla training after SA. For WM, it is indicated that after TNA-PFN, the required iterations are shortened while the interpolated accuracies are similar. The results reveal that training-time neuron alignment can reduce the costs of post-matching and remain similar or even better post-matched LMC.
Table 4: Results of loss barriers on more deep learning tasks.
@@
More deep learning tasks. We conduct more experiments beyond vision tasks and display the results in Table 4. Polynomial approximation task (Peña et al., 2023; von Oswald et al., 2019): we use an MLP with one hidden layer to approximate the second and third polynomial functions: $y=2x^{2}-1,y=(x-3)^{3}$. Sentiment analysis of text (Liu et al., 2022b): we use an LSTM (Graves & Graves, 2012) to predict the sentiment of IMDb reviews (Maas et al., 2011). It can be seen that the loss barriers are decreased by training-time alignment under both polynomial approximation and sentiment analysis tasks. We also implement the experiments on a large-scale dataset, the subset of ImageNet (Deng et al., 2009; tin, Accessed: 2023). The result is shown in Table 10 of Appendix.
@@
Results under non-random initializations. We also examine whether training-time alignment can help when the initializations are not random, which commonly occurs in the pretraining-finetuning paradigm and federated learning. As presented in subsubsection 3.2.3, TNA-PFN is more advantageous to improve LMC under non-random initializations. TNA-PFN nearly clears the barriers under the setting of CNN and CIFAR-10. For the pretrained ResNet18, we observe that TNA-PFN reduces the barrier by a large margin while improving individual accuracies.

### v2

@@ III-C1 Method Formulation
where $\odot$ denotes the element-wise Hadamard product, $\eta$ refers to the learning rate, and $\mathbf{g}_{i}$ is its gradients of the optimizer, such as SGD or Adam. After training for $E$ epochs, we validate the LMC with respect to the loss or accuracy barriers in Definitions III.1 and III.2. The method is notated as TNA-PFN.
Figure 3: A simple demonstration of reducing permutation symmetries via fixed neuron anchors. There are 6 networks (1 original + 5 counterparts) which are functionally identical but with different permutations. The number of permutation symmetries can be reduced by asymmetrically fixing some weights (in red). Though this demonstration presents a simple static network permutation, we will show in the main paper that this kind of method can realize better neuron alignment under training dynamics.
@@
Discussions on differences between TNA-PFN and previous gradient-mask-based methods. Applying gradient masks is discovered in previous gradient compression literature of distribution optimization, but our method is different from the previous works in the aspects as follows. i) Motivation difference: Gradient compression is proposed for communication efficiency of distributed optimization while we study the training-time neuron alignment problem in LMC and model fusion. ii) Implementation difference: Gradient compression uses different random top-k gradient masks at each worker and changes the mask per communication iteration [50, 51]; whereas, TNA-PFN uses the same random gradient masks at each model, fixes the mask, and independently trains the models without any communications; and FedPFN/FedPNU (presented in section V) uses the same masks at each client’s local training and changes the mask per global communication round. iii) Effect difference: Since the masks of workers are different and changing, previous gradient compression methods cannot learn in a consistent subspace of parameters, while we learn in a subspace by the same gradient mask so that some neuron weights are not updated.
III-C2 Theoretical Analysis
We make a theoretical analysis about how TNA-PFN can improve LMC, shown in Theorem III.4 (proof is in the Appendix). The main idea is to treat the linear interpolated landscape of the barrier as a function of parameter $\alpha$, and the connectivity can be depicted by the first and second derivatives of the function.
Theorem III.4.
We define a two-layer neural network with ReLU activation, and the function is $f_{{\bm{v}},{\bm{U}}}({\bm{x}})={\bm{v}}^{\top}\sigma({\bm{U}}{\bm{x}})$ where $\sigma(\cdot)$ is the ReLU activation function. ${\bm{v}}\in\mathbb{R}^{h}$ and ${\bm{U}}\in{\mathbb{R}}^{h\times d}$ are parameters [^2] For simplicity and without loss of generality, the bias terms are omitted. and ${\bm{x}}\in{\mathbb{R}}^{d}$ is the input which is taken from ${\mathbb{X}}=\{{\bm{x}}\in{\mathbb{R}}^{d}|\|{\bm{x}}\|_{2}<b\}$ uniformly. Consider two different networks parameterized with $\{{\bm{U}},{\bm{v}}\}$ and $\{{\bm{U}}^{\prime},{\bm{v}}^{\prime}\}$ respectively, and for arbitrarily chosen masks ${\bm{M}}_{\bm{v}}\in\{0,1\}^{h}$ and ${\bm{M}}_{{\bm{U}}}\in\{0,1\}^{h\times d}$, each element of $\bm{U}$} and $\bm{U}$$^{\prime}$, $\bm{v}$} and $\bm{v}$$^{\prime}$ is i.i.d. sampled from a sub-Gaussian distribution $\mathcal{\text{sub-G}}(0,\sigma_{\bm{U}}^{2})$ and $\mathcal{\text{sub-G}}(0,\sigma_{\bm{v}}^{2})$ respectively with setting ${v}_{i}={v}^{\prime}_{i}$ when ${M}_{{\bm{v}},i}=0$ and ${U}_{i,j}={U}^{\prime}_{i,j}$ when ${M}_{{\bm{U}},ij}=0$. We consider the linear mode connectivity of the two networks and define the difference function between interpolated network and original networks as $z_{{\bm{x}}}(\alpha)=(\alpha{\bm{v}}+(1-\alpha){\bm{v}}^{\prime})^{\top}\sigma((\alpha{\bm{U}}+(1-\alpha){\bm{U}}^{\prime}){\bm{x}})-\alpha{\bm{v}}^{\top}\sigma({\bm{U}}{\bm{x}})-(1-\alpha){{\bm{v}}^{\prime}}^{\top}\sigma({\bm{U}}^{\prime}{\bm{x}})$, $\alpha\in[0,1]$. The function over all inputs is defined as $z(\alpha)=\frac{1}{|{\mathbb{X}}|}\int_{\mathbb{X}}z_{{\bm{x}}}(\alpha)d{\bm{x}}$. We use $\left|z(\alpha)\right|$, $\left|\frac{dz(\alpha)}{d\alpha}\right|$ and $\left|\frac{d^{2}z(\alpha)}{d\alpha^{2}}\right|$ to depict the linear mode connectivity, showing the output changes along the $\alpha$ path. With probability $1-\delta$, it has,
@@ III-C2 Theoretical Analysis
TABLE I: The performances of post-matching methods after TNA-PFN. Interpolated Accuracy (Interp. Acc.) means the accuracy of the linearly interpolated model, i.e., $\mathcal{A}(0.5\mathbf{w}_{1}+0.5\mathbf{w}_{2})$. “Iter.” refers to the number of iterations in the post-matching methods, reflecting the computation costs.
@@
Theorem III.4 shows how TNA-PFN improves LMC by masking some weights from updating (setting $\rho_{\bm{v}}$ and $\rho_{\bm{U}}$ larger than zeros), where the networks learn in a unified permutation subspace.
III-C3 Empirical Results on Linear Mode Connectivity
@@
Permutation symmetry: an intuitive example. To provide neuron-level evidence, we design a demonstration example that showcases and visualizes permutation symmetry during standard training and the permutation asymmetry encouraged by TNA-PFN. We use the synthetic data and MLP in the 2nd Polynomial dataset of Table IV, and we use an initialization of ordered weights for better visualization and validation. The neuron weight changes from initialization to the final model are visualized, and it is found that TNA-PFN can anchor the unfrozen weights. Results are shown in Figures 1 and 2 of the Appendix.
@@
Different model depths, widths, and architectures. In Figure 4, we conduct experiments on MLP with different hidden layers and widths. For MNIST [52], we find shallower and wider networks will not cause barriers, which is consistent with the previous observations [18]. For CIFAR-10 [53], the barriers always exist under various depths and widths. Our proposed TNA-PFN can obviously reduce the accuracy barriers from 0.3-0.4 to 0.1, and we also visualize the loss landscapes, which illustrate the barrier reductions.
We study the LMC of simple CNN and ResNets and present the results in Figure 5. ResNets [54] have higher barriers than simple CNN, and the barriers are exacerbated when the networks are deeper or wider. It is suggested that TNA-PFN can lower the barriers under different architectures.
Generally, we observe that TNA-PFN has more dominant advantages when the models are wider, and the observations are: (1) the second figure in Figure 4: for CIFAR-10 when width increases, the barriers of vanilla training go up while the barriers of TNA-PFN go down; (2) Figure 5: the barrier reduction of TNA-PFN is more obvious for WRN56 compared with ResNet56. This observation is consistent with previous literature about permutation symmetry and linear mode connectivity [18].
TABLE II: Linear mode connectivity of non-random initializations. Dataset: CIFAR-10. Semi-trained CNN: The initialization is first trained on a disjoint subset of CIFAR-10 for 0.5 epoch. Pretrained ResNet18: Pretrained on ImageNet.
@@
TABLE III: Linear mode connectivity of multi-model fusion. The number of models is 5.
@@
The role of post-hoc neuron alignment methods after training-time alignment. We consider simulated annealing (SA) [18] and weight matching (WM) [19] after TNA-PFN in Table I. SA requires large computations, and we notice the improvements are also marginal. Under limited computation budgets (10 or 100 iterations), we find that TNA-PFN can reach a higher result than vanilla training after SA. For WM, it is indicated that after TNA-PFN, the required iterations are shortened while the interpolated accuracies are similar. The results reveal that training-time neuron alignment can reduce the costs of post-matching and remain similar or even better post-matched LMC.
TABLE IV: Results of loss barriers on more deep learning tasks.
@@
More deep learning tasks. We conduct more experiments beyond vision tasks and display the results in Table IV. i) Polynomial approximation task [44, 55]: we use an MLP with one hidden layer to approximate the second and third polynomial functions: $y=2x^{2}-1,y=(x-3)^{3}$. ii) Sentiment analysis of text [56]: we use an LSTM [57] to predict the sentiment of IMDb reviews [58]. It can be seen that the loss barriers are decreased by training-time alignment under both polynomial approximation and sentiment analysis tasks. iii) Large-scale dataset: We also implement the experiments on a large-scale dataset, the subset of ImageNet [59, 60]. The result is shown in the Appendix. In addition, we have conducted the experiments under regression tasks and generative language tasks in Figure 6. We trained MLP regression models with one hidden layer and a width of 16 on the California Housing Dataset [61] and evaluated $R^{2}$ and MAE barriers across runs. We fine-tuned GPT-2 (small) [62] on WikiText-2 [63] under multiple seeds and mild hyperparameter variations. The results in Figure 6 show that TNA-PFN can effectively reduce the connectivity barriers under these tasks.
Figure 6: The connectivity barriers of vanilla training and TNA-PFN under regression and generative tasks. The regression task is training MLP models on the California Housing Dataset with $\rho=0.3$ for TNA-PFN, and the generative task is finetuning GPT-2 (small) on WikiText-2 with $\rho=0.5$ for TNA-PFN.
@@
Results under non-random initializations. We also examine whether training-time alignment can help when the initializations are not random, which commonly occurs in the pretrained-finetuned paradigm and federated learning. As presented in Table III, TNA-PFN is more advantageous to improve LMC under non-random initializations. TNA-PFN nearly clears the barriers under the setting of CNN and CIFAR-10. For the pretrained ResNet18, we observe that TNA-PFN reduces the barrier by a large margin while improving individual accuracies.